OpenStudy (anonymous):

Simplify 6(k – 3) – 5(k + 2).

OpenStudy (anonymous):

6(k - 3) - 5(k + 2) Multiply 6 and 5 by everything inside the parenthesis they are with. 6k - 18 - 5k + 12 Combine like terms. k - 6

OpenStudy (anonymous):

6k-18-5k-10 k-28

OpenStudy (anonymous):

@eclipse18 um... -5 times 2 does not equal 12

OpenStudy (anonymous):

6(k-3)-5(k+2)=0 distribute 6 and -5 6k-18-5k-10=0 k-28=0 k=28

OpenStudy (anonymous):

@XoOo13 K DOES NOT EQUAL 28

OpenStudy (kuoministers):

you guys... it says simplify so simply it means k-28

OpenStudy (whpalmer4):

@eclipse18 you incorrectly turned \(-5(k+2) \) into \(-5k + 10\) because you forgot to apply the negative sign when multiplying -5 and 2. It's a common mistake!

OpenStudy (whpalmer4):

Actually, you turned it into -5k + 12, I can't explain how you got the 12, but 10 I can :-)

OpenStudy (anonymous):

Yes that what I had said the answer was K=-28 6(k – 3) – 5(k + 2) because when you takr 6 times the K you get 6k and than take the six times -3 and get 6k-18 than take the 5 times the K you get -5k than take -5 times +2 and -5k-10 next you take 6k-5k=k than you -18-10 because we -.- we add the -18-10 which results in K=-28

OpenStudy (anonymous):

@whpalmer4 Yeah I realized that later, sorry. I was tired and accidently did -6 times 2, not -5 times 2. My bad.